Consider an infinite surface charge in y-z plane with density σ. Electric field at any point in positive x-axis is σ/2ϵ (i hat) and electric field at any point in negative x-axis is -σ/2ϵ (i hat). Thus we see there is a discontinuity of σ/ϵ at the surface (at x=0). What will be the electric field on the surface charge? That is, what will be the electric field at x=0?
1
Expert's answer
2019-04-29T09:32:30-0400

What will be the electric field on the surface charge?


"E=\frac {\sigma} {2ɛ}"

what will be the electric field at x=0?


"E=\frac {\sigma} {2ɛ} - \frac {\sigma} {2ɛ}=0"